OpenStudy (anonymous):

(Functions) Find f(-5) and f(2) for the given function. f(x)=|x-1| I forgot how to do it.. can someone explain please?

OpenStudy (tkhunny):

Have you considered substituting? f(x) = |x-1| f(3) = |3-1| f(12) = |12-1| \(f(\sqrt{\pi}) = |\sqrt{\pi}-1|\) f(Plastic) = |Plastic-1|

OpenStudy (anonymous):

what is the absolute value of a number?

OpenStudy (tkhunny):

|3| = 3 |-3| = 3